# Network Alignment with Transferable Graph Autoencoders

### v3

@@ Appendix B Spectral characterization of GNNs
where $\lambda_{n},\bm{v}_{n}$ are the $n-$th eigenvalue and eigenvector and $a_{n}^{(i)}=\bm{v}_{n}^{T}\bm{X}^{(l)}\sum_{k=0}^{K-1}\lambda_{n}^{k}\bm{H}_{k}^{(l)}[:,i]$ is a scalar related to the Graph Fourier Transform (GFT) of $\bm{X}^{(l)}$ [Sardellitti et al., 2017]. It is clear from equation (11) that the output of each layer is a linear combination of the adjacency eigenvectors, followed by a pointwise non-linearity. Thus, a GNN can produce unique and more powerful graph embeddings than spectral methods by processing the eigenvectors and eigenvalues of the adjacency matrix.
Appendix C Proof of Theorem 3.2
To prove Theorem 3.2. We consider one layer GNN with a vector input $\bm{x}\in\mathbb{R}^{N}$. This GNN can be represented by the following equation:
@@
C.1 White random input and variance computation
@@
$\mathbb{E}\left[\bm{y}_{i}\right]$ $=\mathbb{E}\left[\text{diag}\left(\sum_{k=0}^{K-1}h_{k}^{i}\bm{S}^{k}\bm{x}\bm{x}^{T}\sum_{j=0}^{K-1}h_{j}^{i}\bm{S}^{j^{T}}\right)\right]=\text{diag}\left(\sum_{k=0}^{K-1}h_{k}^{i}\bm{S}^{k}\mathbb{E}\left[\bm{x}\bm{x}^{T}\right]\sum_{j=0}^{K-1}h_{j}^{i}\bm{S}^{j^{T}}\right)$
@@
C.2 Single band filtering
@@
$\bm{y}=\sum_{k=0}^{K-1}h_{k}\bm{S}^{k}\bm{x}=\sum_{k=0}^{K-1}h_{k}\bm{V}\bm{\Lambda}^{k}\bm{V}^{T}\bm{x}=\sum_{k=0}^{K-1}h_{k}\sum_{n=1}^{N}\lambda_{n}^{k}\bm{v}_{n}\bm{v}_{n}^{T}\bm{x}=\sum_{n=1}^{N}\bm{v}_{n}^{T}\bm{x}\sum_{k=0}^{K-1}h_{k}\lambda_{n}^{k}\bm{v}_{n}.$ (16)
@@ C.3 GNN and absolute eigenvectors
and then equation (C.1) takes the form:
$\mathbb{E}\left[\bm{y}_{i}\right]=\text{diag}\left(\sum_{k=0}^{K-1}h_{k}^{i}\bm{S}^{k}\sum_{j=0}^{K-1}h_{j}^{i}\bm{S}^{j^{T}}\right)=\text{diag}\left(\bm{v}_{\lambda}\bm{v}_{\lambda}^{T}\bm{v}_{\lambda}\bm{v}_{\lambda}^{T}\right)=\text{diag}\left(\bm{v}_{\lambda}\bm{v}_{\lambda}^{T}\right)=\left|\bm{v}_{\lambda}\right|^{2}$ (22)
@@
Appendix D Data Augmentation-Perturbation Model
Figure 4 gives an illustrative example of data augmentation process, used for training, and the perturbation model used during testing. In particular, we add or remove an edge with a certain probability yielding ${\bm{S}^{\prime}}=\bm{S}+\bm{M}$, where $\bm{M}\in\{-1,0,1\}^{N\times N}$. Note that $\bm{M}[i,j]$ can be equal to $1$ and $-1$ only if $\bm{S}[i,j]$ is equal to 00 and $1$ respectively. Then we relabel each node by permuting the rows and columns of the adjacency matrix yielding $\hat{\bm{S}}=\bm{P}\left(\bm{S}+\bm{M}\right)\bm{P}^{T}$.
Figure 4: The proposed data augmentation process. The first step is to remove existing edges and/or add new ones, while the second step reorders the nodes by permuting the rows and columns of the graph adjacency.
@@ Appendix E Implementation Details
In this section we discuss the implementation details of our framework. The code and datasets used in all experiments can be found here: https://github.com/GraphMatching/Network-Alignment-with-Transferable-Graph-Autoencoders
@@ E.1 Assignment Optimization
The problem in (28) can be optimally solved in $\mathcal{O}\left(N^{3}\right)$ flops by the Hungarian algorithm [Kuhn, 1955a]. To avoid this computational burden we employ the greedy Hungarian approach that has computational complexity $\mathcal{O}\left(N^{2}\right)$ and usually works well in practice.
@@ E.2 Datasets
The descriptions of the real-world datasets used in our experiments are presented below:
@@ E.3.2 Graph/Node embedding techniques
  • DeepWalk [Perozzi et al., 2014]: A node embedding approach, simulates random walks on the graph and apply skip-gram on the walks to generate node embedding. We use the implementation from Karateclub https://github.com/benedekrozemberczki/karateclub/blob/master/karateclub/node_embedding/neighbourhood/deepwalk.pyThe algorithm is implemented with the default parameters as suggested by this repository, the number of random walks is 10 with each walk of length 80. The dimensionality of embedding is set to be 128. We run the algorithm with 1 epoch and set the learning rate to be 0.05.
  • Node2Vec [Grover and Leskovec, 2016b]: An improve version of DeepWalk, it has weights on the randomly generated random walks, to make the neighborhood preserving objective more flexible. We use the implementation from https://github.com/benedekrozemberczki/karateclub/blob/master/karateclub/node_embedding/neighbourhood/node2vec.py. The default parameters are used. We simulate 10 random walks on the graph with length 80. p and q are both equal to 1. Dimensionality of embeddings is set to be 4 and we run 1 epoch with learning rate 0.05.
  • GraphWave [Donnat et al., 2018]: The structure information of the graphs is captured by simulating heat diffusion process on them. We use the implementation from [Donnat et al., 2018]:https://github.com/benedekrozemberczki/karateclub/blob/master/karateclub/node_embedding/structural/graphwave.py with the default parameters: number of evaluation points is 200, step size is 0.1, heat coefficient is 1.0 and Chebyshev polynomial order is set to be 100. Note that this implementation does not work on graphs with more than 10,000 nodes, so we exclude this model on the DBLP and Coauthor_CS dataset.
@@ E.3.3 Optimization based graph matching algorithms
  • S-GWL [Xu et al., 2019a] matches two given graphs by retrieving node correspondence from the optimal transport associated with the Gromov-Wasserstein discrepancy between the graphs. We use the implementation by the authors in https://github.com/HongtengXu/s-gwl, since the performance of S-GWL is very sensitive to the parameter gamma, as suggested by the authors, we fine-tuned this parameter over the range of $[0.001,0.1]$ for each dataset on the cleaned graph, and use that optimal parameter for all other experiments on this dataset.
E.4 T-GAE Parameters
We implement our proposed Transferable graph autoencoder (T-GAE) framework, following Equations (6), (7), (8), for the tasks described later in the section.
For the graph matching experiments in Tables 3 and 5. We use 1 MLP and 8 GNN layers, followed by another MLP to design the Encoder $\phi\left(\bm{X},{\bm{S}};\mathcal{H}\right)$ of T-GAE, as illustrated in Fig. 1. For the GNN layers we use graph isomorphism network (GIN) [Xu et al., 2019b], which is described by (4) with $f$ being the summation function, and $g$ being the MLP. For both graph matching and sub-graph matching experiments, we use hidden dimension of 16 and output dimension of 8. To ensure the fairness of comparison, this set of parameters are tuned on matching the cleaned version of Celegans, no other fine-tuning is conduced for each individual dataset or perturbation levels.
For the sub-graph matching experiments in Table 4, on ACMDBLP, we use 1 MLP and 12 GNN layers, followed by another MLP as encoder, the hidden and output dimension is set to 1024. On Douban OnlineOffline, we use 6 GNN layers with hidden and output dimensions equal to 512. This set of parameters is chosen based on the best matching accuracy achieved.
To summarize, detailed settings of our experiments are:
  • Optimizer, epoch and learning rate: We train our model using Adam optimizer for 200 epochs and initial learning rate equal to $0.001$ for all graph matching experiments. For sub-graph matching experiments, we train the model using Adam optimizer for 50 epochs and the initial learning rate is set to $0.0001$.
  • Dimensions: The hidden dimension of the GNN modules is equal to $16$ and that of the output layer (encoder) is equal to 8 for all graph matching experiments. For sub-graph matching experiments, on ACMDBLP, the hidden and output dimension equal to 1024. On Douban OnlineOffline, they are both set to 512.
  • Number of layers: On all graph matching experiments, the number of GNN layers equals to 8, and for sub-graph matching experiments, we use 12 GNN layers for ACMDBLP and 6 GNN layers for Douban OnlineOffline.
  • Activation function: The activation function for each GNN layer is ReLU.
  • Decoder: We use an 1-layer decoder followed by the sigmoid function.
  • Data Augmentation: To conduct the experiments in Tables 3 and 5, we augment the dataset with $10$ randomly generated perturbed graphs for each perturbation level and perturbation model.
  • Compute Resources: All our experiments are conduced on AMD Ryzen Threadripper 3960X 24-Core Processor and Nvidia RTX A6000 GPU.
Appendix F Efficiency and Run-time Comparison
We analyze the efficiency of the proposed graph matching framework by comparing its running time with the competing algorithms. The GNN-based methods achieve at most $\times$50 less running time on both graph matching and subgraph matching tasks. The proposed training objective (8) scales well from networks with 1,000 nodes [Zhang and Tong, 2016] to denser networks with $\times$10 nodes, with minor running time increase, compared to other node embedding techniques.
We also noticed that compared to the other GNN based framework WAlign Gao et al. [2021], T-GAE takes more time to train on the larger scale networks as illustrated in Figure 5. This is because the deep GNN is trained by the structure of the network. The training takes more time as the netowrk scales up and the number of nodes and edges grow dramatically, whereas the training labels for WAlign come from the input feature, which does not scale up with as the size of the training graphs increase. This could be one of the limitations to the proposed training paradigm and that is why we leverage transfer learning to efficiently match large scale graphs, as illustrated in Table3.
Table 2: Runtime in seconds for the competing algorithms on graph matching tasks.
Algorithm Celegans Arenas Douban Cora Dblp Coauthor CS
T-GAE 0.0980.098{0.098}0.098 0.2410.241{0.241}0.241 1.9341.934{1.934}1.934 0.8420.842{0.842}0.842 173.836173.836{173.836}173.836 193.289193.289{193.289}193.289
GAE 0.0940.094{0.094}0.094 0.2520.252{0.252}0.252 1.9491.949{1.949}1.949 0.8570.857{0.857}0.857 174.410174.410{174.410}174.410 194.062194.062{194.062}194.062
WAlign 0.0780.078{0.078}0.078 0.2050.205{0.205}0.205 1.8001.800{1.800}1.800 0.7660.766{0.766}0.766 169.694169.694{169.694}169.694 189.032189.032{189.032}189.032
Spectrul 5.7125.712{5.712}5.712 2.8192.819{2.819}2.819 60.29860.298{60.298}60.298 54.77054.770{54.770}54.770 > 48 hours > 48 hours
GraphWave 8.3088.308{8.308}8.308 32.99432.994{32.994}32.994 281.629281.629{281.629}281.629 131.230131.230{131.230}131.230 5470.7245470.724{5470.724}5470.724 6291.3686291.368{6291.368}6291.368
ConeAlign 1.3331.333{1.333}1.333 3.5003.500{3.500}3.500 31.95531.955{31.955}31.955 13.79913.799{13.799}13.799 887.090887.090{887.090}887.090 1099.1451099.145{1099.145}1099.145
S-GWL 27.84427.844{27.844}27.844 37.44337.443{37.443}37.443 3394.5223394.522{3394.522}3394.522 311.201311.201{311.201}311.201 > 48 hours > 48 hours
FINAL 0.0300.030{0.030}0.030 0.0810.081{0.081}0.081 1.0071.007{1.007}1.007 0.4980.498{0.498}0.498 86.78886.788{86.788}86.788 118.065118.065{118.065}118.065
Figure 5: Runtime comparison for the competing algorithms on the sub-graph matching datasets. All GNN-based methods have 8 layers with 512 as hidden dimensions and are trained with 20 epochs.
Appendix G More Baseline Results
@@
Appendix H Degree Perturbation Model Results
@@
$5\%$ perturbation Celegans ${14.4\pm 3.9}$ ${12.6\pm 1.4}$ ${2.6\pm 0.6}$ ${24.3\pm 2.3}$ ${39.1\pm 21.1}$ $\mathbf{60.3\pm 1.9}$ ${6.8\pm 0.9}$ ${2.0\pm 0.7}$ ${18.3\pm 1.4}$
Arenas ${28.3\pm 10.0}$ ${22.4\pm 1.0}$ ${0.5\pm 0.4}$ ${20.5\pm 1.6}$ $\mathbf{92.2\pm 0.9}$ ${82.2\pm 1.2}$ ${6.9\pm 1.1}$ ${0.0\pm 0.0}$ ${60.1\pm 1.9}$
Douban ${1.4\pm 0.7}$ ${10.3\pm 0.4}$ ${0.3\pm 0.0}$ ${7.5\pm 0.4}$ $\mathbf{56.1\pm 12.3}$ ${48.2\pm 2.2}$ ${5.2\pm 0.9}$ ${0.1\pm 0.0}$ ${29.2\pm 2.3}$
Cora ${3.5\pm 1.9}$ ${32.0\pm 0.7}$ ${0.1\pm 0.0}$ ${8.6\pm 0.7}$ ${15.5\pm 5.5}$ ${24.7\pm 3.0}$ ${24.2\pm 1.4}$ ${4.0\pm 0.5}$ $\mathbf{48.5\pm 1.2}$
Dblp ${0.3\pm 0.1}$ ${18.4\pm 0.1}$ doesn’t scale ${4.7\pm 0.2}$ > 48 hours ${10.0\pm 0.8}$ ${14.9\pm 0.2}$ ${5.0\pm 0.3}$ $\mathbf{36.4\pm 1.2}$
Coauthor CS ${0.5\pm 0.2}$ ${20.2\pm 0.3}$ doesn’t scale ${3.5\pm 1.3}$ > 48 hours $\mathbf{47.7\pm 2.7}$ ${4.6\pm 0.1}$ ${0.5\pm 0.1}$ ${35.3\pm 0.5}$
@@
It is also noticed that when we introduce $5\%$ perturbation, the performance of all competing methods dropped compared to the results in Table 3. This is because the Uniform model is agnostic to the graph structure. In the Degree perturbation model however, an edge with higher degree is more likely to be removed, so it changes the structure of the graph more drastically, which makes the task of graph matching more challenging.

### v4

@@ D.1 Spectral characterization of GNNs
where $\lambda_{n},\bm{v}_{n}$ are the $n-$th eigenvalue and eigenvector and $a_{n}^{(i)}=\bm{v}_{n}^{T}\bm{X}^{(l)}\sum_{k=0}^{K-1}\lambda_{n}^{k}\bm{H}_{k}^{(l)}[:,i]$ is a scalar related to the Graph Fourier Transform (GFT) of $\bm{X}^{(l)}$ [Sardellitti et al., 2017]. It is clear from equation (10) that the output of each layer is a linear combination of the adjacency eigenvectors, followed by a pointwise non-linearity. Thus, a GNN can produce unique and more powerful graph embeddings than spectral methods by processing the eigenvectors and eigenvalues of the adjacency matrix. To prove Theorem 3.2. We consider one layer GNN with a vector input $\bm{x}\in\mathbb{R}^{N}$. This GNN can be represented by the following equation:
@@
D.2 White random input and variance computation
@@
$\mathbb{E}\left[\bm{y}_{i}\right]$ $=\mathbb{E}\left[\text{diag}\left(\sum_{k=0}^{K-1}h_{k}^{i}\bm{S}^{k}\bm{x}\bm{x}^{T}\sum_{j=0}^{K-1}h_{j}^{i}\bm{S}^{j^{T}}\right)\right]$ (14)
$=\text{diag}\left(\sum_{k=0}^{K-1}h_{k}^{i}\bm{S}^{k}\mathbb{E}\left[\bm{x}\bm{x}^{T}\right]\sum_{j=0}^{K-1}h_{j}^{i}\bm{S}^{j^{T}}\right)$
@@
D.3 Single band filtering
@@
$\bm{y}$ $=\sum_{k=0}^{K-1}h_{k}\bm{S}^{k}\bm{x}$ (16)
$=\sum_{k=0}^{K-1}h_{k}\bm{V}\bm{\Lambda}^{k}\bm{V}^{T}\bm{x}$ (17)
$=\sum_{k=0}^{K-1}h_{k}\sum_{n=1}^{N}\lambda_{n}^{k}\bm{v}_{n}\bm{v}_{n}^{T}\bm{x}$ (18)
$=\sum_{n=1}^{N}\bm{v}_{n}^{T}\bm{x}\sum_{k=0}^{K-1}h_{k}\lambda_{n}^{k}\bm{v}_{n}.$ (19)
@@ D.4 GNN and absolute eigenvectors
and then equation (14) takes the form:
$\mathbb{E}\left[\bm{y}_{i}\right]$ $=\text{diag}\left(\sum_{k=0}^{K-1}h_{k}^{i}\bm{S}^{k}\sum_{j=0}^{K-1}h_{j}^{i}\bm{S}^{j^{T}}\right)$ (25)
$=\text{diag}\left(\bm{v}_{\lambda}\bm{v}_{\lambda}^{T}\bm{v}_{\lambda}\bm{v}_{\lambda}^{T}\right)$ (26)
$=\text{diag}\left(\bm{v}_{\lambda}\bm{v}_{\lambda}^{T}\right)$ (27)
$=\left|\bm{v}_{\lambda}\right|^{2}$ (28)
@@
D.5 Generalization to multiple graph pairs
The analysis in the previoius subsections is indeed presented for a pair of graphs but can be directly extended for any set of graphs. We can generalize the Theorem 3.2, to read as: Let $\{\mathcal{G}_{1},\dots,\mathcal{G}_{M}\}$ be a set of graphs with adjacencies $\{\bm{S}_{1},\dots,\bm{S}_{M}\}$ that have non-repeated eigenvalues. Then for any $\bm{S},~{}\hat{\bm{S}}\in\{\bm{S}_{1},\dots,\bm{S}_{M}\}$, there exists a GNN $\phi\left(\bm{X};\bm{S},\mathcal{H}\right):\mathbb{R}^{N\times D}\to\mathbb{R}^{N\times D^{L}}$ such that:
with
$\bm{P}^{\ast}=\operatorname*{arg\,min}_{\bm{P}\in\mathcal{P}}~{}\left\lVert~{}\phi\left(\bm{X};{\bm{S}},\mathcal{H}\right)-\bm{P}\phi\left(\hat{\bm{X}};\hat{\bm{S}},\mathcal{H}\right)~{}\right\rVert_{F}^{2},$
where ${\bm{P}^{\diamond}},~{}\check{\bm{P}}$ are solutions to the optimization problems in (1) and (3) respectively.
@@ Appendix E Implementation Details
In this section we discuss the implementation details of our framework.
@@ E.1 Assignment Optimization
The problem in (34) can be optimally solved in $\mathcal{O}\left(N^{3}\right)$ flops by the Hungarian algorithm [Kuhn, 1955b]. To avoid this computational burden we employ the greedy Hungarian approach that has computational complexity $\mathcal{O}\left(N^{2}\right)$ and usually works well in practice.
@@ E.2 Datasets
Table 2: Summary of Dataset statistics that are included in Section 5
Graph Matching Celegans [Kunegis, 2013] 453 2,025 2,025 Interactome
Arenas [Leskovec and Krevl, 2014] 1,133 5,451 5,451 Email Communication
Cora [Sen et al., 2008] 2,708 5,278 5,278 Citation Network
Douban [Zhang and Tong, 2016] 3,906 7,215 7,215 Social Network
Dblp [Pan et al., 2016] 17,716 52,867 52,867 Citation Network
Coauthor CS [Shchur et al., 2018] 18,333 81,894 81,894 Coauthor Network
Subraph Matching ACM-DBLP [Zhang and Tong, 2019] 9,872 9,916 39,561 44,808 6,352 Citation Network
Douban Online-Offline [Zhang and Tong, 2016] 3,906 1,118 1,632 3,022 1,118 Social Network
We include statistics of the datasets used in our experiments in Table 2. The detailed descriptions of each dataset are presented below:
@@ E.3.2 Graph/Node embedding techniques
  • DeepWalk [Perozzi et al., 2014]: A node embedding approach, simulates random walks on the graph and apply skip-gram on the walks to generate node embedding. We use the implementation from Karateclub. The algorithm is implemented with the default parameters as suggested by this repository, the number of random walks is 10 with each walk of length 80. The dimensionality of embedding is set to be 128. We run the algorithm with 1 epoch and set the learning rate to be 0.05.
  • Node2Vec [Grover and Leskovec, 2016b]: An improve version of DeepWalk, it has weights on the randomly generated random walks, to make the neighborhood preserving objective more flexible. We use the implementation from Karateclub. The default parameters are used. We simulate 10 random walks on the graph with length 80. p and q are both equal to 1. Dimensionality of embeddings is set to be 4 and we run 1 epoch with learning rate 0.05.
  • GraphWave [Donnat et al., 2018]: The structure information of the graphs is captured by simulating heat diffusion process on them. We use the implementation from Karateclub with the default parameters: number of evaluation points is 200, step size is 0.1, heat coefficient is 1.0 and Chebyshev polynomial order is set to be 100. Note that this implementation does not work on graphs with more than 10,000 nodes, so we exclude this model on the DBLP and Coauthor_CS dataset.
@@ E.3.3 Optimization based graph matching algorithms
  • S-GWL [Xu et al., 2019b] matches two given graphs by retrieving node correspondence from the optimal transport associated with the Gromov-Wasserstein discrepancy between the graphs. We use the implementation by the authors in https://github.com/HongtengXu/s-gwl, since the performance of S-GWL is very sensitive to the parameter gamma, as suggested by the authors, we fine-tuned this parameter over the range of $[0.001,0.1]$ for each dataset on the cleaned graph, and use that optimal parameter for all other experiments on this dataset.
E.4 T-GAE model details
As illustrated in Figure 1, the structure of our proposed encoder consists of two MLPs and a series of GNN layers. The node features are processed by a 2-layer MLP and passed to all the GNN layers. We add skip connections between this MLP layer and all the subsequent GNN layers. The outputs of all GNN layers are concatenated and passed to another 2-layer MLP, followed by a linear decoder to generate the reconstructed graph. The model is optimized end to end by equation 7. For graph matching experiments, since we consider the general case where graphs are given without node attributes, we use the $7$ structural features proposed in [Berlingerio et al., 2013]. The features include the degree of each node, the local and average clustering coefficient, and the number of edges, outgoing edges, and neighbors in each node’s egonet. This input feature is applied for all GNN-based methods. As a result, the performance of NetSimile, vanilla GAE and WAlign provide measures to assess the benefit of using T-GAE for node embedding. Note that one can choose different message passing functions as $f$ and $g$ in Equation (4), and any structure-preserving node features. Our reported results are based on GIN [Xu et al., 2019a] and Netsimile [Berlingerio et al., 2013].
Appendix F More Baseline Results
@@
Appendix G Degree Perturbation Model Results
@@
Appendix H Efficiency analysis
H.1 Running time comparison
T-GAE is a scalable and efficient approach for network alignment. We analyze the efficiency of the proposed graph matching framework by comparing its running time with the competing algorithms. T-GAE achieves at most $\times$2000 less running time, on graph matching tasks, compared to the optimization-based methods, as shown in Table 6. Compared to the existing GNN based approaches, T-GAE consistently achieves the shortest training time and inference time, this is because we replace some message passing layers by the local MLP which serves as an attention function on the output of all GNN layers, such models are empirically proved to be more efficient, at the same time, prompting the expressiveness of GNN layers, to generate node embedding for graph matching. It should be noted that T-GAE is also the only approach that is transferable, which means it does not need to re-train on every pair of new graphs. T-GAE greatly improves the scalability of optimization based methods, as well as the effectiveness of the existing GNN frameworks.
[…]
The proposed training objective (7) scales well from networks with 400 nodes [Kunegis, 2013] to denser networks with $\times$50 nodes [Shchur et al., 2018], with minor running time increase, compared to other GNN-based frameworks(WAlign), as shown in Figure 5.
Table 6: Runtime(inference+matching) in seconds for the competing algorithms on graph matching tasks.
Algorithm Celegans Arenas Cora Douban Dblp Coauthor CS
Spectrul 5.7125.712{5.712}5.712 2.8192.819{2.819}2.819 54.77054.770{54.770}54.770 60.29860.298{60.298}60.298 > 48 hours > 48 hours
Netsimile 1.2121.212{1.212}1.212 1.5601.560{1.560}1.560 1.6161.616{1.616}1.616 4.4004.400{4.400}4.400 195.542195.542{195.542}195.542 225.546225.546{225.546}225.546
GraphWave 8.3088.308{8.308}8.308 32.99432.994{32.994}32.994 131.230131.230{131.230}131.230 281.629281.629{281.629}281.629 5470.7245470.724{5470.724}5470.724 6291.3686291.368{6291.368}6291.368
FINAL(Matlab) 0.0300.030{0.030}0.030 0.0810.081{0.081}0.081 0.4980.498{0.498}0.498 1.0071.007{1.007}1.007 86.78886.788{86.788}86.788 118.065118.065{118.065}118.065
S-GWL 27.84427.844{27.844}27.844 37.44337.443{37.443}37.443 311.201311.201{311.201}311.201 3394.5223394.522{3394.522}3394.522 > 48 hours > 48 hours
ConeAlign 1.3331.333{1.333}1.333 3.5003.500{3.500}3.500 13.79913.799{13.799}13.799 31.95531.955{31.955}31.955 887.090887.090{887.090}887.090 1099.1451099.145{1099.145}1099.145
WAlign 0.0780.078{0.078}0.078 0.2050.205{0.205}0.205 0.7660.766{0.766}0.766 1.8001.800{1.800}1.800 169.694169.694{169.694}169.694 189.032189.032{189.032}189.032
GAE 0.0740.074{0.074}0.074 0.2120.212{0.212}0.212 0.7570.757{0.757}0.757 1.7491.749{1.749}1.749 164.410164.410{164.410}164.410 184.062184.062{184.062}184.062
T-GAE(ours) $\mathbf{0.068}$ $\mathbf{0.201}$ $\mathbf{0.742}$ $\mathbf{1.734}$ $\mathbf{163.836}$ $\mathbf{183.289}$
H.2 Matching algorithms comparison
In this subsection, we evaluate the accuracy and matching time of different matching algorithms, to demonstrate how matching algorithms of different time complexity influence the performance of the proposed T-GAE framework. To guarantee fairness of comparison, we use an untrained T-GAE encoder to encode the graphs, and use (1) approximated NN algorithm introduced in Section 4.3 of time complexity $\mathcal{O}(NlogN)$. (2) greedy Hungarian algorithm as applied in Section 5 and 5.3 of time complexity $\mathcal{O}(N^{2})$. (3) exact Hungarian algorithm [Kuhn, 1955b] of time complexity $\mathcal{O}(N^{3})$, where $N$ is the number of nodes in the graph. We report average accuracy and running time on matching 10 randomly generated samples.
Table 7: Graph matching performance and matching time on Celegans and Arenas using untrained T-GAE encoder. We report results of approximated NN matching algorithm, greedy Hungarian algorithm, and exact Hungarian algorithm. We highlight the performance gain of exact Hungarian over the approximated NN and Greedy Hungarian.
# Dataset/Perturbation Celegans Arenas
0 0.01 0.05 0 0.01 0.05
Acc
Time Acc Time Acc Time Acc Time Acc Time Acc Time
$\mathcal{O}(NlogN)$
1 Approximated NN $89.8\pm 1.0$ 0.004 $0.8\pm 0.3$ 0.004 $0.7\pm 0.2$ 0.004 $98.0\pm 2.6$ 0.009 $0.1\pm 0.1$ 0.009 $0.0\pm 0$ 0.0 0.009
$\mathcal{O}(N^{2})$
2 Greedy Hungarian $88.4\pm 0.9$ 0.068 $80.3\pm 0.3$ 0.068 $58.2\pm 3.5$ 0.067 $97.6\pm 0.4$ 0.173 $93.1\pm 0.5$ 0.176 $60.2\pm 5.2$ 0.174
$\mathcal{O}(N^{3})$
3 Exact Hungarian $88.5\pm 0.8$ 0.225 $84.0\pm 0.2$ 24.778 $64.7\pm 1.8$ 73.674 $97.6\pm 0.4$ 0.411 $93.8\pm 0.4$ 562.221 $70.0\pm 2.0$ 1624.713
$\mathcal{O}(N^{3})$
4 Acc Gain in % (Exact vs Approx/Greedy) -1.3 / +0.1 +83.2 / +3.7 +64.0 / +6.5 -0.4 / 0 +93.7 / +0.7 +70.0 / +9.8
The performance of T-GAE to match small graphs can be further improved by adopting the exact Hungarian algorithm. Comparing Exact Hungarian and Greedy Hungarian, we observe that when there is no perturbation involved, greedy Hungarian and exact Hungarian achieves comparable performance. However, the exact Hungarian algorithm outperforms the greedy version in occurrence of perturbations, and the performance gap increases as we introduce more topology noise. Specifically, it offers a 6.5% and 9.8% matching accuracy increase on Celegans and Arenas, respectively, on an untrained T-GAE encoder. However, it can take more than $\times$9000 longer than the greedy algorithm, on Arenas 5% perturbation, for example. This implies that whenever applicable, the exact Hungarian algorithm should be applied to further improve the performance of T-GAE to match small graphs, especially when the noise level is high, but there is a trade-off between matching accuracy and efficiency.
The efficiency of T-GAE to match permuted graphs can be enhanced by applying the approximated NN algorithm. We observe that the approximated NN algorithm we introduced in Section 4.3 effectively match the aligned nodes of permuted graphs, and on Arenas dataset of 1,133 nodes, it saves 95% matching time compared to the greedy Hungarian algorithm. However, this algorithm fails to match perturbed graphs. This is because the 1-dimensional feature is not expressive enough to to catch the topological noise. Our experiments prove that this efficient NN algorithm should be applied when we match very large scale networks with their permuted versions.
Overall, we divide graph matching using T-GAE in three scenarios: (1) Matching small graphs, exact Hungarian algorithm should be applied. (2) Matching large scale networks without perturbation, the approximation NN algorithm should be deployed to enhance efficiency. (3) The greedy Hungarian algorithm provides a good trade-off between time and efficiency for the general form of graph matching.